### v3

# arXiv:2602.05968v3[math.AP] 28 Feb 2026

Abstract.
In this paper, we obtain stability results for the $L^{p}$-Poincaré inequality for both Lebesgue and Gaussian probability measures (Theorem 3.3 and Theorem 3.13) that involve explicit dependence on the geometry of the domain. As a byproduct, the explicit constant allows us to recover important results of Yu, Zhong [YZ86] and Smits [SMI96] (Corollary 3.9), related to the fundamental gap conjecture of the Laplacian (resolved by Andrews and Clutterbuck [AC11]), thereby providing an alternative proof. Moreover, we extend this spectral gap result to the $p$-Laplacian (Corollary 3.6). Such gap estimates for the Dirichlet $p$-Laplacian appear to be unavailable, as also observed in [DSW18]. Our approach relies on properties of the first eigenfunction of the (Gaussian) $p$-Laplacian operator and weighted Poincaré inequalities for log-concave measures on convex domains.
Key words and phrases:
Poincaré inequality, stability inequalities, Gaussian Poincaré inequality, spectral gap
2020 Mathematics Subject Classification:
26D10, 35J60, 60E15
This research is funded by the Committee of Science of the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP23490970).
1. Introduction
Suppose that there is a functional inequality $A(u)\geq B(u)$ for all $u$ in some space of functions. Let $\mathcal{M}$ denote the space of all functions which achieve an equality. The basic question here is the following: Is the inequality sensitive to small changes? In the functional sense, that means if $A(u)\geq B(u)$ is very close to an equality, does that imply that $u$ is close to the set of optimizers $\mathcal{M}$? The positive answer is usually given in the form of an inequality
$A(u)-B(u)\geq cd(u,\mathcal{M}),$ (1.1)
where $c>0$ and $d(u,\mathcal{M})$ is the appropriate distance function to the set of optimizers $\mathcal{M}$.
This question was first raised by Brezis and Lieb in [BL85] for the $L^{2}$-Sobolev inequality and later resolved completely by Bianchi and Egnell in [BE91]. In particular, Bianchi and Egnell obtained that for all $u\in W_{0}^{1,2}(\mathbb{R}^{n})$, the following stability inequality holds:
$\int_{\mathbb{R}^{n}}|\nabla u|^{2}dx-S_{n}\left(\int_{\mathbb{R}^{n}}|u|^{\frac{2n}{n-2}}dx\right)^{\frac{n-2}{n}}\geq c_{S}\inf_{U\in E_{Sob}}\int_{\mathbb{R}^{n}}|\nabla(u-U)|^{2}dx,$
where $c_{S}>0$, $S_{n}$ is the sharp Sobolev constant and $E_{Sob}$ is the manifold of optimizers of the Sobolev inequality. The literature on the stability analysis of Sobolev inequality is extensive and for the interested reader we refer to [BWW03, CIA06, FMP07, CFM+09, FMP13, FUS15, FN19, NEU20, FZ22a, DEF+25].
Following the contributions of Brezis, Lieb, Bianchi and Egnell, stability results have been investigated for various inequalities, including isoperimetric [FMP08, FMP10, CL12, BP12, FMM18, FZ22b], Brunn-Minkowski [FMP09, FJ17, vST21, vST19, FvT23, FvT24], Gagliardo-Nirenberg [RUF14, DT16, NGU19, BDN+25, ZZ25, CT25b], Hardy [CF08, MOW13, MOW15, SAN18, RSY18, RS18, RSY25, BGR26], Heisenberg-Pauli-Weyl [MV21, FAT21, CFL+24a, HY25, DGL+24, SYZ25, DLL26] and $L^{2}$-Poincaré-Wirtinger [LLR26] with Gaussian weights. We also refer to [FIG13] for an interesting survey with applications to the long-time asymptotic of evolution equations. In this work, however, we are specifically focused on the stability of the $L^{2}$-Poincaré inequality.
Let $1<p<\infty$ and $\Omega$ be an open bounded domain of $\mathbb{R}^{n}$. Then, for every $u\in W_{0}^{1,p}(\Omega)$, we have
$\int_{\Omega}|\nabla u|^{p}dx\geq\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx,$ (1.2)
where $\lambda_{1}(p,\Omega)$ is the first eigenvalue of the Dirichlet $p$-Laplacian. The literature on the stability analysis of the inequality (1.2) seems to be very limited. However, there are some works with remainder terms [FT02, BK23]. For example, in [FT02], Fleckinger-Pellé and Takáč obtained an improvement of the inequality (1.2): Assume that $\Omega$ is either an interval in $\mathbb{R}^{1}$ or else a bounded domain in $\mathbb{R}^{n}$ ($n\geq 2$) whose boundary $\partial\Omega$ is a compact connected $C^{2}$-manifold. Then there exists a constant $c\equiv c(p,\Omega)>0$ such that for all $u\in W_{0}^{1,p}(\Omega)$
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}\int_{\Omega}|u|^{p}dx\geq c\left(|u^{\parallel}|^{p-2}\int_{\Omega}|\nabla u_{1}|^{p-2}|\nabla u^{\perp}|^{2}dx+\int_{\Omega}|\nabla u^{\perp}|^{p}dx\right),$ (1.3)
where
$u=u^{\parallel}u_{1}+u^{\perp},\quad u^{\parallel}:=\|u_{1}\|_{L^{2}(\Omega)}^{-2}\langle u,u_{1}\rangle\in\mathbb{R}\quad\text{and}\quad\langle u^{\perp},u_{1}\rangle=0$ (1.4)
and $\langle\cdot,\cdot\rangle$ is the usual scalar product in $L^{2}$. It is important to note that the result (1.3), in [FT02], was then used to obtain an existence result concerning a (p1)(p-1)-homogeneous problem with Dirichlet $p$-Laplacian.
Then, Bobkov and Kolonitskii [BK23] later extended (1.3) to a Friedrichs inequality with a slightly modified version of the orthogonal decomposition than in (1.4). Furthermore, leveraging the results in [DS04, BL23], Bobkov and Kolonitskii were able to relax the regularity assumptions on $\Omega$ and confirm the conjecture proposed in [PS04, Section 2.1] concerning the relationship between the hypotheses in [FT02]. At this point, again, we note that the stability analysis was not performed in [BK23] and is similar in structure to [FT02] as they then also apply the corresponding result to a similar boundary value problem at resonance as an application of the improved Friedrichs inequality.
In a very recent paper [LLR26], Lam, Lu and Russanov obtained a certain version of the gradient stability of the Poincaré-Wirtinger with Gaussian probability measure in $L^{2}(\mathbb{R}^{n})$. They subsequently applied this result to derive an improved Poincaré-Wirtinger inequality with Gaussian weights, which they then utilized to investigate the stability of the related Heisenberg-Pauli-Weyl (HPW) inequality. These results have also been generalized to monomial weights in the same paper [LLR26].
The idea of connecting the HPW identity with a suitable Poincaré inequality to establish stability results for the HPW or Caffarelli-Kohn-Nirenberg inequalities has been successfully employed before (see, e.g. [DFL+23, CFL+24a, DGL+24, SYZ25, DLL26]). In this paper, we show how such approach can also be adapted to obtain stability results for the Poincaré inequalities. We do this by utilizing some of the properties of the first eigenfunction of the $p$-Laplacian [SAK87, CQS25] and the weighted Poincaré inequality for log-concave measures [FNT12]. As a result, we are able to derive the following stability result with explicit constant: Let $p\geq 2$, $\Omega$ be an open bounded convex set having diameter $\text{diam}(\Omega)$ with smooth boundary $\partial\Omega$. Then, for every $u\in C_{0}^{\infty}(\Omega)$, we have
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}d(u,E_{Poin})^{p}.$ (1.5)
Here, $E_{Poin}=\{cu_{1}:c\in\mathbb{R}\}$ is the manifold of optimizers (or eigenspace) of (1.2),
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}$
and $d(u,E_{Poin})$ is the $L^{p}$ distance to the manifold of optimizers, i.e.
$d(u,E_{Poin})^{p}:=\inf_{c\in\mathbb{R}}\left\{\left\lVert u-cu_{1}\right\rVert^{p}_{L^{p}(\mathbb{R}^{n})}\right\}.$
The constant, in (1.5), is explicit with clear dependence on the diameter of the convex domain $\Omega$, which we believe is the first such result. In general, in most stability results, the explicit form of the constant that appears nearby the distance function (1.1) is unknown. Even for the historic result of Bianchi and Egnell [BE91] for the Sobolev inequality, which was obtained in 1991, no information was known for the constant $c_{S}$ (except an upper bound [KÖN23]) until very recently [DEF+25]. There, Dolbeault, Esteban, Figalli, Frank and Loss obtained sharp stability results for the Sobolev inequality with optimal dimensional dependence (see also, [CLT+25]). In our case, the inequality (3.2) provides not only stability for the $L^{p}$-Poincaré inequality, but also gives an explicit geometric dependence of the stability constant due to the result of Ferone, Nitsch and Trombetti [FNT12]. Additionally, we show some corresponding results for the Gaussian probability measure using the same logic thanks to [CQS25].
Moreover, after normalizing the eigenfunctions $u_{1}$ and $u_{2}$, we obtain the following spectral gap of the Dirichlet $p$-Laplacian:
$\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}C(p,\Omega,u_{1},u_{2}),$ (1.6)
where
$C(p,\Omega,u_{1},u_{2}):=\inf_{c\in\mathbb{R}}\int_{\Omega}|u_{2}-cu_{1}|^{p}dx.$
We could not find any results related to the spectral gaps for the Dirichlet $p$-Laplacian operator. In fact, in [DSW18, Page 7], Dai, Seto and Wei, note that “similar gap estimate for the $p$-Laplacian is still unknown”. We believe (1.6) is the first such estimate.
Specializing to $p=2$ in (1.6), we get the fundamental gap of the Dirichlet Laplacian in the form of Yu, Zhong and Smits [YZ86, SMI96]:
$\lambda_{2}(2,\Omega)-\lambda_{1}(2,\Omega)\geq\frac{\pi^{2}}{\text{diam}(\Omega)^{2}}.$
The difference between the first two eigenvalues of the Dirichlet Laplacian on a bounded open domain $\Omega$, denoted by $\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)$, is commonly known as the fundamental gap of $\Omega$. The fundamental gap problem has an interesting history and we give a brief overview of it here.
Originally, van den Berg, in [VAN83], first observed that many convex sets have a spectral gap of $\frac{3\pi^{2}}{\text{diam}(\Omega)^{2}}$. This was also independently observed and then conjectured to be true in any convex domain $\Omega$ by Yau [YAU87], Ashbaugh and Benguria [AB89] for the Schrödinger operator $\Delta+V$. The general statement in terms of the Schrödinger operator has been resolved in the one-dimensional setting by Lavine [LAV94] (see, also, [AB89, HOR03]). In higher dimensions, a breakthrough was made by Singer, Wong, Yau and Yau [SWY+85], where they obtained a lower bound of $\frac{\pi^{2}}{4\text{diam}(\Omega)^{2}}$ by utilizing the gradient estimates in spirit of those from Li [LI79], Li and Yau [LY80] and the log-concavity of the first eigenfunction (established for convex potentials by Brascamp and Lieb [BL76]). Later, Yu, Zhong [YZ86] and [SMI96] improved the estimate to $\frac{\pi^{2}}{\text{diam}(\Omega)^{2}}$. Under certain assumptions on the domain $\Omega$, major step forward in proving the conjecture was made by Bañuelos and Méndez-Hernández [BM00], Davis [DAV01] and Bañuelos with Kröger [BK01]. Finally, Andrews and Clutterbuck [AC11] proved the conjecture in 2011 for any convex domain $\Omega$. A few years later, an alternative proof was given by Ni in [NI13]. Gap estimates on other domains are addressed in [CO97, OSW99, SWW19, HWZ20, DSW21]. In a very recent paper [ABF24], Amato, Bucur and Fragalà strengthened the result of Andrews and Clutterbuck and quantified it in terms of flatness, thereby answering an additional open problem by Yau [YAU93]. Additional works related to the fundamental gap problem of the Neumann Laplacian are present in [PW60, ZHO84]. We also refer to [AND14, DSW18] for comprehensive surveys on the subject.
2. Preliminaries
In this section, we briefly recall the necessary notation and provide some preliminary results. Let $x\in\mathbb{R}^{n}$ be a point in the $n$-dimensional Euclidean space $\mathbb{R}^{n}$. Then, $|x|$ denotes the Euclidean norm of $x$ in $\mathbb{R}^{n}$ with $dx$ denoting the integration with respect to the standard Lebesgue measure.
Let $1<p<\infty$, $\Omega$ be a bounded domain of $\mathbb{R}^{n}$. Then, a non-trivial function $u\in W_{0}^{1,p}(\Omega)$ is called an eigenfunction of the $p$-Laplacian $\Delta_{p}=\text{div}\left(|\nabla u|^{p-2}\nabla u\right)$ with a corresponding eigenvalue $\lambda\in\mathbb{R}$ if it satisfies the following in the weak sense:
$\begin{cases}-\Delta_{p}u=\lambda|u|^{p-2}u&\text{in }\Omega,\\ u=0&\text{on }\partial\Omega.\end{cases}$ (2.1)
When $p=2$, the problem (2.1) reduces to the eigenvalue problem for the standard Laplace operator. In general, the first eigenvalue can be characterized as
$\lambda_{1}(p,\Omega)=\inf\left\{\frac{\int_{\Omega}|\nabla u|^{p}dx}{\int_{\Omega}|u|^{p}dx}:u\in W_{0}^{1,p}(\Omega)\setminus\{0\}\right\}.$ (2.2)
We remark that $\lambda_{1}(p,\Omega)>0$ and is attained by the associated first eigenfunction $u_{1}\in W_{0}^{1,p}(\Omega)$. All first eigenfunctions are known to coincide up to modulo scaling (see, [VÁZ84, LIN90, BK02]) and do not change sign in $\Omega$. Therefore, it is possible to assume without the loss of generality that $u_{1}>0$ unless stated otherwise. Furthermore, $u_{1}$ is known to be bounded and in $C_{loc}^{1,\beta}(\Omega)$ for some $\beta\in(0,1)$ (for more details, see, [DIB83, ÔTA88, TOL84]).
When $p\neq 2$, one can construct several infinite sequences of variational eigenvalues by the Ljusternik-Schnirelman theory with topological indexes such as Krasnosel’skii genus $\gamma(\mathcal{A})$ of a symmetric set $\mathcal{A}\subset W_{0}^{1,p}(\Omega)$ (see, [GP87]). In particular, let $\Sigma_{k}$ be the collection of all symmetric subsets $\mathcal{A}$ of $W_{0}^{1,p}(\Omega)$ with $\gamma(\mathcal{A})\geq k$. Then, the numbers
$\lambda_{k}(p,\Omega)=\inf_{\mathcal{A}\in\Sigma_{k}}\sup_{u\in\mathcal{A}}\frac{\int_{\Omega}|\nabla u|^{p}dx}{\int_{\Omega}|u|^{p}dx}$ (2.3)
form a sequence of eigenvalues such that
$0<\lambda_{1}(p,\Omega)<\lambda_{2}(p,\Omega)\leq\cdots\leq\lambda_{k}(p,\Omega)\to+\infty\quad\text{as }k\to+\infty.$
When $p=2$, the eigenvalues of (2.3) coincide with the standard discrete sequence of eigenvalues of the Laplacian (see, [BPS16, ABM14]). Unlike the $p=2$ case, it is unknown whether these are all eigenvalues of the $p$-Laplacian. It is known, however, that there are no eigenvalues between $\lambda_{1}(p,\Omega)$ and $\lambda_{2}(p,\Omega)$ with $\lambda_{1}(p,\Omega)$ being the smallest eigenvalue (see, [AT96]). Then, Cuesta, De Figueiredo and Gossez, in [CDG99], derived a variational characterization of $\lambda_{2}(p,\Omega)$ for an open bounded connected domain $\Omega$ (later proved for any open domain with finite measure by Brasco and Franzina [BF13]):
$\lambda_{2}(p,\Omega)=\inf_{\gamma\in\Gamma(u_{1},-u_{1})}\left[\max_{u\in\gamma([0,1])}\int_{\Omega}|\nabla u|^{p}dx\right],$ (2.4)
where $u_{1}$ is the first eigenfunction such that $\left\lVert u_{1}\right\rVert_{L^{p}(\Omega)}=1$ and $\Gamma(u_{1},-u_{1})$ is the family of all continuous maps from $[0,1]$ to $\mathcal{A}_{p}(\Omega)=\{u\in W_{0}^{1,p}(\Omega):\|u\|_{L^{p}(\Omega)}=1\}$ with endpoints $u_{1}$ and $u_{1}$. We refer also to [DR99, PER03, JL05, BD17, ABP18, FMZ19, BT25, BĞ25, BOB26] for other results in this direction.
We finish this part by stating an important result of Sakaguchi [SAK87, Theorem 1] on the log-concavity of the eigenfunctions of the Dirichlet $p$-Laplacian operator (see, also, [BL76]).
Theorem 2.1 ([SAK87]).
Let $\Omega$ be a bounded convex domain in $\mathbb{R}^{n}\ (n\geq 2)$ with smooth boundary $\partial\Omega$. Fix a number $p>1$. Let $u\in W_{0}^{1,p}(\Omega)$ be a positive weak solution to the nonlinear eigenvalue problem
$\begin{cases}-\Delta_{p}u=\lambda|u|^{p-2}u&\text{in }\Omega,\\ u=0&\text{on }\partial\Omega.\end{cases}$
Then, $v=\log u$ is a concave function.
On the other hand, we have the Gaussian probability space $(\mathbb{R}^{n},\gamma)$ with measure $\gamma$ defined as
$\gamma(\Omega)=(2\pi)^{-\frac{n}{2}}\int_{\Omega}e^{-\frac{|x|^{2}}{2}}dx$
for any measurable set $\Omega\subseteq\mathbb{R}^{n}$. In this context, we denote integration with respect to the measure $\gamma$ as $d\gamma$. This measure is explicitly defined by the density $d\gamma(x)=(2\pi)^{-\frac{n}{2}}e^{-\frac{|x|^{2}}{2}}\,dx$. Furthermore, $d\gamma_{\partial\Omega}$ represents the (n1)(n-1)-dimensional Hausdorff measure on the boundary $\partial\Omega$ with respect to $\gamma$.
Now let $1<p<\infty$ and $\Omega\subseteq\mathbb{R}^{n}$ be a measurable set, then the meaning of the notation $L^{p}(\Omega)$, $W^{1,p}(\Omega)$ and $W_{0}^{1,p}(\Omega)$ is same as usual. In the setting of the Gaussian probability measure, we define $L^{p}(\Omega,\gamma)$ as the space of all measurable functions $u$ such that
$L^{p}(\Omega,\gamma)=\left\{u:\Omega\to\mathbb{R}:\|u\|_{p,\gamma}^{p}:=\int_{\Omega}|u(x)|^{p}d\gamma<+\infty\right\}.$
In the same spirit, we have $W^{1,p}(\Omega,\gamma)$, $W_{0}^{1,p}(\Omega,\gamma)$ as the $\gamma$-weighted Sobolev spaces. We note that if $\Omega$ is a bounded set, then the density $e^{-\frac{|x|^{2}}{2}}$ has both positive upper and lower bounds in $\Omega$, thereby implying $L^{p}(\Omega,\gamma)=L^{p}(\Omega)$, $W^{1,p}(\Omega,\gamma)=W^{1,p}(\Omega)$ and etc. Otherwise, when $\Omega$ is unbounded, we have $L^{p}(\Omega)\subset L^{p}(\Omega,\gamma)$ and $W^{1,p}(\Omega)\subset W^{1,p}(\Omega,\gamma)$ (see, e.g., [KIL94, HKM18, PRR19]). In addition, we define the notion of the Gaussian divergence:
$\text{div}_{\gamma}X:=e^{\frac{|x|^{2}}{2}}\text{div}\left(e^{-\frac{|x|^{2}}{2}}X\right)=\text{div}X-\langle X,x\rangle$
along with the corresponding product rule:
$\text{div}_{\gamma}\left(fV\right)=\langle\nabla f,V\rangle+f\text{div}_{\gamma}V.$
Let $\Omega$ be a bounded domain of $\mathbb{R}^{n}$. Similarly as for the Euclidean $p$-Laplacian, we define the Gaussian $p$-Laplacian operator as
$-\Delta_{p,\gamma}u:=-\operatorname{div}(|\nabla u|^{p-2}\nabla u)+\langle x,\nabla u\rangle|\nabla u|^{p-2},$
which is an extension of the Ornstein-Uhlenbeck operator as for $p=2$, we recover its usual form
$L_{\gamma}u:=-\Delta u+\langle x,\nabla u\rangle.$
Under appropriate regularity of functions and domains, the integration by parts formula is as follow:
$\int_{\Omega}v\Delta_{p,\gamma}(u)d\gamma=-\int_{\Omega}|\nabla u|^{p-2}\langle\nabla u,\nabla v\rangle d\gamma+\int_{\partial\Omega}v|\nabla u|^{p-2}\langle\nabla u,n_{x}\rangle d\gamma_{\partial\Omega},$
where $n_{x}$ is the outward unit normal vector at point $x\in\partial\Omega$. We call $u\in W_{0}^{1,p}(\Omega,\gamma)$ an eigenfunction of the $p$-Gaussian operator associated to the eigenvalue $\lambda\in\mathbb{R}$ if it is a weak solution to
$\begin{cases}-\Delta_{p,\gamma}u=\lambda|u|^{p-2}u&\text{ in }\Omega,\\ u=0&\text{ on }\partial\Omega.\end{cases}$ (2.5)
For the existence of solutions to (2.5) we refer to [FPS+23, Section 6]. Here, we define its first positive eigenvalue
$\lambda_{1}(p,\Omega,\gamma)=\inf\left\{\frac{\int_{\Omega}|\nabla u|^{p}d\gamma}{\int_{\Omega}|u|^{p}d\gamma}:u\in W_{0}^{1,p}(\Omega,\gamma)\setminus\{0\}\right\}.$
Just as in the Euclidean case, without the loss of generality, one can take the first eigenfunction $u^{\gamma}_{1}$ of the Gaussian $p$-Laplacian as positive as it does not change sign in $\Omega$ and is unique up to scalar multiplication (for more details, see, [DMW+21, FPS+23]).
Finally, we recall a corresponding result of Sakaguchi for the Gaussian $p$-Laplacian operator by Colesanti, Qin, Salani [CQS25, Theorem 1.5], some regularity properties of the Gaussian eigenfunctions [CQS25, Proposition 2.1], the weighted $L^{p}$-Poincaré inequality for the log-concave probability measures (see, [MAZ13, Section 6.4.3] and [FNT12]) as well as sharp bounds on the constant [FNT12, Theorem 1.1]:
Theorem 2.2 ([CQS25, Theorem 1.5]).
Let $p>1$, $n\geq 2$, $\Omega$ be an open, bounded and convex domain in $\mathbb{R}^{n}$, and $u$ be a solution of problem (2.5), with $u>0$ in $\Omega$. Then the function
$W=\log u$
is concave in $\Omega$.
We remark that, in Theorem 2.2, Colesanti, Qin and Salani do not require the boundary $\partial\Omega$ to be smooth compared to Theorem 2.1. So, it might be possible to remove the smoothness condition for the Euclidean $p$-Laplacian case too.
Proposition 2.3 ([CQS25, Proposition 2.1]).
If $\Omega$ is a bounded open domain and $u$ is a weak solution of problem (2.5), with $p>1$, then $u\in C^{1,\alpha}_{\text{loc}}(\Omega)\cap C^{2}(\Omega\setminus\bar{\mathcal{C}})$ for some $\alpha\in(0,1)$, where $\mathcal{C}=\{x\in\Omega:\nabla u(x)=0\}$. Moreover, if $\partial\Omega$ is $C^{1,\alpha}$, then $u\in C^{1,\beta}(\bar{\Omega})$, for some $\beta\in(0,1)$.
Theorem 2.4 ([FNT12, Theorem 1.1]).
Let $p>1$ and $\omega$ be a positive log-concave function on an open bounded convex set $\Omega$ having the diameter $\text{diam}(\Omega)$. Then, there exists a positive constant $C_{\Omega,p,\omega}$ such that, for every Lipschitz function $u$ with $\int_{\Omega}|u|^{p-2}u\omega dx=0$, we have
$\int_{\Omega}\left|\nabla u\right|^{p}\omega dx\geq C_{\Omega,p,\omega}\inf_{t\in\mathbb{R}}\int_{\Omega}\left|u-t\right|^{p}\omega dx.$
Moreover, in any dimension, we have
$C_{\Omega,p,\omega}\geq\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p},$
where
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}.$
For the explicit expression of $\pi_{p}$, we refer to [STA90, LIN95, STA93, RW99, BBČ+06]. In the next part of this section, we introduce the definition of the $C_{p}$-functional and its upper and lower bounds, which will be essential for our proof.
Definition 2.5.
Let $1<p<\infty$. Then, for $\xi,\eta\in\mathbb{C}^{n}$, we define
$C_{p}(\xi,\eta):=|\xi|^{p}-|\xi-\eta|^{p}-p|\xi-\eta|^{p-2}\langle\textnormal{Re}(\xi-\eta),\overline{\eta}\rangle\geq 0.$ (2.6)
Lemma 2.6 ([CKL+24b, Step 3 of Proof of Theorem 1.2]).
Let $p\geq 2$. Then, for $\xi,\eta\in\mathbb{C}^{n}$, we have
$C_{p}(\xi,\eta)\geq c_{1}(p)|\eta|^{p},$ (2.7)
where
$c_{1}(p)=\inf_{(s,t)\in\mathbb{R}^{2}\setminus\{(0,0)\}}\frac{\bigl[t^{2}+s^{2}+2s+1\bigr]^{\frac{p}{2}}-1-ps}{\bigl[t^{2}+s^{2}\bigr]^{\frac{p}{2}}}\in(0,1].$
The constant, in (2.7), is especially interesting. In [CKL+24b], the authors asked whether it is the most optimal for the given inequality. This question was recently answered by Huang and Tong [HT25] and we state it here.
Lemma 2.7 ([HT25, Theorem 6]).
Let $p\geq 2$. The sharp constant $c_{1}(p)$ in (2.7) is determined as
$c_{1}(p)=(p-1)(1-k_{0})^{p}+pk_{0}(1-k_{0})^{p-1}+k_{0}^{p}>0,$ (2.8)
where $k_{0}=\frac{r_{0}}{1+r_{0}}$ and $r_{0}$ is the solution of the equation
$r^{p-1}-(p-1)r-(p-2)=0.$ (2.9)
Example 2.8.
For $p=3$, the sharp constant is given by $c_{1}(3)=2-\sqrt{2}$.
Remark 2.9.
Note that since $k_{0}=\frac{r_{0}}{1+r_{0}}$, we have $1-k_{0}=\frac{1}{1+r_{0}}$. Substituting this into (2.8) gives us
$c_{1}(p)=(p-1)\left(\frac{1}{1+r_{0}}\right)^{p}+p\frac{r_{0}}{1+r_{0}}\left(\frac{1}{1+r_{0}}\right)^{p-1}+\left(\frac{r_{0}}{1+r_{0}}\right)^{p}.$
This simplifies to
$c_{1}(p)=\frac{p-1+pr_{0}+r^{p}_{0}}{(1+r_{0})^{p}}.$ (2.10)
From equation (2.9), we get that
$r^{p}_{0}=r^{2}_{0}(p-1)+r_{0}(p-2).$ (2.11)
Now we substitute (2.11) to (2.10):
$c_{1}(p)$ $=\frac{p-1+pr_{0}+r^{2}_{0}(p-1)+r_{0}(p-2)}{(1+r_{0})^{p}}$
$=\frac{p-1+(2p-2)r_{0}+r^{2}_{0}(p-1)}{(1+r_{0})^{p}}$
$=(p-1)(r_{0}+1)^{2-p}.$
This is a much simpler expression for $c_{1}(p)$ and it is the one which we will use throughout the text.
Although Lemma 2.7 makes it possible to compute the sharp constant $c_{1}(p)$ via a simple polynomial equation, it is unclear what happens as $p\to\infty$, which is interesting on its own right. That is why, using Lemma 2.7, we provide lower and upper bounds for the constant $c_{1}(p)$ (for more details see Section 4) that exactly reveal its behavior as $p\to\infty$.
Proposition 2.10.
For any $p\geq 2$, we have
$\frac{1}{2^{p-2}}\leq c_{1}(p)\leq\frac{p-1}{2^{p-2}}.$
Moreover, as $p\to\infty$, the constant $c_{1}(p)\to 0$.
Lemma 2.11 ([CT25a, Lemma 2.2]).
Let $1<p<2\leq n$. Then, for $\xi,\eta\in\mathbb{C}^{n}$, we have
$C_{p}(\xi,\eta)\geq c_{2}(p)\frac{|\eta|^{2}}{\left(|\xi|+|\xi-\eta|\right)^{2-p}},$
where
$c_{2}(p):=\inf_{s^{2}+t^{2}>0}\frac{\left(t^{2}+s^{2}+2s+1\right)^{\frac{p}{2}}-1-ps}{\left(\sqrt{t^{2}+s^{2}+2s+1}+1\right)^{p-2}(t^{2}+s^{2})}\in\left(0,\frac{p(p-1)}{2^{p-1}}\right].$
Lemma 2.12 ([CT25a, Lemma 2.3]).
Let $1<p<2\leq n$. Then, for $\xi,\eta\in\mathbb{C}^{n}$, we have
$C_{p}(\xi,\eta)\leq c_{3}(p)\frac{|\eta|^{2}}{\left(|\xi|+|\xi-\eta|\right)^{2-p}},$
where
$c_{3}(p):=\sup_{s^{2}+t^{2}>0}\frac{\left(t^{2}+s^{2}+2s+1\right)^{\frac{p}{2}}-1-ps}{\left(\sqrt{t^{2}+s^{2}+2s+1}+1\right)^{p-2}(t^{2}+s^{2})}\in\left[\frac{p}{2^{p-1}},+\infty\right).$
Lastly, we use the Picone-type identity that connects the $C_{p}$-functional (2.6) from [AYZ25, Theorem 3.13]:
Theorem 2.13 ([AYZ25]).
Let $u$ be a complex-valued function on $\Omega\subset\mathbb{R}^{n}$ and $\phi$ be a nontrivial complex-valued function on $\Omega\subset\mathbb{R}^{n}$. Then, we have
$C_{p}(\xi,\eta)=|\nabla u|^{p}+(p-1)\left|\frac{\nabla\phi}{\phi}u\right|^{p}-p\,\textnormal{Re}\left[\left|\frac{\nabla\phi}{\phi}u\right|^{p-2}\frac{u}{\phi}\langle\nabla\phi,\overline{\nabla u}\rangle\right],$
$R_{p}(\xi,\eta)=|\nabla u|^{p}-|\nabla\phi|^{p-2}\biggl\langle\nabla\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right),\nabla\phi\biggr\rangle$
and
$C_{p}(\xi,\eta)=R_{p}(\xi,\eta)\geq 0,$
where $C_{p}(\cdot,\cdot)$ is given in (2.6) and
$\xi:=\nabla u,\quad\eta:=\nabla u-\frac{\nabla\phi}{\phi}u.$
We refer to the crucial work of Allegretto and Huang [AX98], where the Picone’s identity for real-valued functions made its first appearance.
3. Main results
In this section, we state the stability of the $L^{p}$-Poincaré inequality. First, we show results concerning the usual Euclidean $p$-Laplacian, then for the Gaussian $p$-Laplacian. We then apply the stability inequality to establish known and new fundamental gap estimates.
Theorem 3.1 ([AYZ25]).
Let $1<p<\infty$, $\Omega\subset\mathbb{R}^{n}$ be a set where the divergence theorem holds. Then, for all complex-valued $u\in C^{\infty}_{0}(\Omega)$, we have
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx=\int_{\Omega}C_{p}\left(\nabla u,u_{1}\nabla\left(\frac{u}{u_{1}}\right)\right)dx.$ (3.1)
Moreover, the remainder term vanishes if and only if $u=cu_{1}$ for all $c\in\mathbb{C}$.
Remark 3.2.
The identity, as with all such identities, thanks to Lemmata 2.6, 2.11 and 2.12 allows us to derive all possible optimizers of the $L^{p}$-Poincaré inequality for any $p\in\left(1,\infty\right)$. Such set, we will denote as $E_{Poin}:=\{cu_{1}:c\in\mathbb{C}\}$. If we restrict the $L^{p}$-Poincaré inequality to real-valued functions, then the set of all optimizers becomes $E_{Poin}=\{cu_{1}:c\in\mathbb{R}\}$.
The identity (3.1) is the starting point in proving the stability result. In combination with the Theorem 2.1 and Theorem 2.4, we finally obtain the stability result of the $L^{p}$-Poincaré inequality:
Theorem 3.3.
Let $p\geq 2$, $\Omega$ be an open bounded convex set having diameter $\text{diam}(\Omega)$ with smooth boundary $\partial\Omega$. Then, for every $u\in C_{0}^{\infty}(\Omega)$, we have
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}d(u,E_{Poin})^{p}.$ (3.2)
Here, $E_{Poin}=\{cu_{1}:c\in\mathbb{R}\}$ is the manifold of optimizers (or eigenspace) of (1.2),
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}$
and $d(u,E_{Poin})$ is the $L^{p}$ distance to the manifold of optimizers, i.e.
$d(u,E_{Poin})^{p}:=\inf_{c\in\mathbb{R}}\left\{\left\lVert u-cu_{1}\right\rVert^{p}_{L^{p}(\mathbb{R}^{n})}\right\}.$
When $p=2$, in (3.2), interestingly the constant $\pi_{2}=\pi$, which implies:
Corollary 3.4.
Let $\Omega$ be an open bounded convex set having diameter $d$ with smooth boundary $\partial\Omega$. Then, for every $u\in C_{0}^{\infty}(\Omega)$, we have
$\int_{\Omega}|\nabla u|^{2}dx-\lambda_{1}(2,\Omega)\int_{\Omega}|u|^{2}dx\geq\left(\frac{\pi}{\text{diam}(\Omega)}\right)^{2}d(u,E_{Poin})^{2}.$ (3.3)
Remark 3.5.
The density argument can be applied to extend the inequalities (3.2) and (3.3) to $u\in W_{0}^{1,p}(\Omega)$, which we will use for obtaining the fundamental gap of the $p$-Laplacian.
The inequality (3.2) actually allows us to obtain a fundamental gap of the Dirichlet $p$-Laplacian on bounded convex domains with smooth boundary:
Corollary 3.6.
Let $p$ and $\Omega$ be from Theorem 3.3 with normalized eigenfunctions, i.e. $\left\lVert u_{1}\right\rVert_{L^{p}(\Omega)}=\left\lVert u_{2}\right\rVert_{L^{p}(\Omega)}=1$. Then, for $\lambda_{1}(p,\Omega)$ and $\lambda_{2}(p,\Omega)$ defined as in (2.2)-(2.4), we have
$\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}C(p,\Omega,u_{1},u_{2}),$ (3.4)
where
$C(p,\Omega,u_{1},u_{2}):=\inf_{c\in\mathbb{R}}\int_{\Omega}|u_{2}-cu_{1}|^{p}dx.$
Remark 3.7.
In the aforementioned survey [DSW18, Page 7], Dai, Seto and Wei comment that spectral gaps for the Dirichlet $p$-Laplacian operator are “still unknown” and we believe it is the first such result in this direction.
Remark 3.8.
It is standard to assume the minimization problem (2.2) on the set $S_{p}=\{u\in W_{0}^{1,p}(\Omega)\ |\left\lVert u\right\rVert_{L^{p}(\Omega)}=1\}$ (see, for example, [LIN90, KL06, ABP18, dRS19]).
When $p=2$, in the inequality (3.4), we have that $C(2,\Omega,u_{1},u_{2})=1$ and $\pi_{2}=\pi$, which gives us the fundamental gap of the Laplacian in the form of Yu, Zhong and Smits [YZ86, SMI96].
Corollary 3.9.
Let $\Omega$ be from Theorem 3.3 with normalized eigenfunctions. Then, for $\lambda_{1}(2,\Omega)$ and $\lambda_{2}(2,\Omega)$ defined as in (2.2)-(2.4), we have
$\lambda_{2}(2,\Omega)-\lambda_{1}(2,\Omega)\geq\frac{\pi^{2}}{\text{diam}(\Omega)^{2}}.$
Remark 3.10.
We remark that, as pointed out to us by V. Bobkov, from the inequality (1.3), the stability of the $L^{p}$-Poincaré inequality in the form of Bianchi and Egnell can be obtained as follows:
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}\int_{\Omega}|u|^{p}dx\geq c\int_{\Omega}|\nabla u^{\perp}|^{p}dx=c\int_{\Omega}\left|\nabla\left(u-u^{\parallel}u_{1}\right)\right|^{p}dx\\ \geq c\inf_{U\in E_{Poin}}\int_{\Omega}\left|\nabla\left(u-U\right)\right|^{p}dx,$ (3.5)
where $E_{Poin}$ is the eigenspace. However, the constant, in (3.5), is not explicit. Here, in this paper, we will demonstrate a different approach, which allows us to obtain a stability result with an explicit constant.
Next, we show some corresponding results for the Gaussian probability measure.
Theorem 3.11.
Let $1<p<\infty$, $\Omega\subset\mathbb{R}^{n}$ be a set where the divergence theorem holds. Then, for all complex-valued $u\in C^{\infty}_{0}(\Omega,\gamma)$, we have
$\int_{\Omega}|\nabla u|^{p}d\gamma-\lambda_{1}(p,\Omega,\gamma)\int_{\Omega}|u|^{p}d\gamma=\int_{\Omega}C_{p}\left(\nabla u,u^{\gamma}_{1}\nabla\left(\frac{u}{u^{\gamma}_{1}}\right)\right)d\gamma.$ (3.6)
Moreover, the remainder term vanishes if and only if $u=cu^{\gamma}_{1}$ for all $c\in\mathbb{C}$.
Remark 3.12.
The identity (3.6) implies that all optimizers of the inequality
$\int_{\Omega}|\nabla u|^{p}d\gamma\geq\lambda_{1}(p,\Omega,\gamma)\int_{\Omega}|u|^{p}d\gamma$
are in the set that we denote $E^{\gamma}_{Poin}:=\{cu^{\gamma}_{1}:c\in\mathbb{R}\ (\mathbb{C})\}$.
By the same argument as in the proof of Theorem 3.3, we obtain the stability of the $L^{p}$-Poincaré inequality for Gaussian probability measures.
Theorem 3.13.
Let $p\geq 2$ and $\Omega$ be an open bounded convex set having the diameter $d$. Then, for every $u\in C_{0}^{\infty}(\Omega,\gamma)$, we have
$\int_{\Omega}|\nabla u|^{p}d\gamma-\lambda_{1}(p,\Omega,\gamma)\int_{\Omega}|u|^{p}d\gamma\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}d(u,E_{Poin}^{\gamma})^{p},$ (3.7)
where
$\pi_{p}=2\int_{0}^{+\infty}\frac{1}{1+\frac{1}{p-1}s^{p}}ds=2\pi\frac{(p-1)^{1/p}}{p(\sin(\pi/p))}$
and $d(u,E^{\gamma}_{Poin})$ is the $L^{p}$ distance to the manifold of optimizers, i.e.
$d(u,E^{\gamma}_{Poin})^{p}:=\inf_{c\in\mathbb{R}}\left\{\left\lVert u-cu^{\gamma}_{1}\right\rVert^{p}_{L^{p}(\mathbb{R}^{n},\gamma)}\right\}.$
Remark 3.14.
As we noted in Section 2, Theorem 2.4 of Colesanti, Qin and Salani does not require the boundary of the domain $\partial\Omega$ to be smooth. Thus, in the case of Gaussian probability measure, Theorem 3.13 does not assume $\partial\Omega$ to be smooth.
In the same spirit as in Corollary 3.4, for $p=2$ in (3.7), we have
Corollary 3.15.
Let $\Omega$ be an open bounded convex set having the diameter $d$. Then, for every $u\in C_{0}^{\infty}(\Omega)$, we have
$\int_{\Omega}|\nabla u|^{2}d\gamma-\lambda_{1}(2,\Omega,\gamma)\int_{\Omega}|u|^{2}d\gamma\geq\left(\frac{\pi}{\text{diam}(\Omega)}\right)^{2}d(u,E^{\gamma}_{Poin})^{2}.$
4. Proof of Proposition 2.10
Proof of Proposition 2.10.
Let us recall that
$c_{1}(p)=(p-1)(r_{0}+1)^{2-p},$ (4.1)
where $r_{0}$ is the root of the function
$f(r)=r^{p-1}-(p-1)r-(p-2).$
When $p=2$, we have $c_{1}(2)=1$. When $p>2$, we first take the derivative
$f^{\prime}(r)=(p-1)(r^{p-2}-1).$
For any $p>2$, we see that $f^{\prime}(r)=0$ at $r=1$. In addition, for any $p>2$ and $r>1$, one has $f^{\prime}(r)>0$. Also note that f(1)=42pf(1)=4-2p starts at a negative value and becomes positive for large $r$. By continuity of $f(r)$, there exists at least one root $r_{0}>1$. Now since $f(r)$ is strictly increasing on $r\in(1,\infty)$, by the Intermediate Value Theorem, the root is unique and satisfies $r_{0}>1$.
Since $r_{0}>1$, we have $r_{0}+1>2$ and consequently $(r_{0}+1)^{p-2}>2^{p-2}$. Therefore, from (4.1), we obtain that for $p>2$:
$c_{1}(p)<\frac{p-1}{2^{p-2}}.$
Combining $p=2$ and $p>2$ cases, we get that
$c_{1}(p)\leq\frac{p-1}{2^{p-2}}.$
To get a lower bound, we construct a new function $g(t)=t^{q}$, which is convex for any $q\geq 1$ and $t>0$. Since $g(t)$ is convex, it is also midpoint convex:
$g\left(\frac{r_{0}+1}{2}\right)\leq\frac{g(r_{0})+g(1)}{2},$
which is equivalent to
$\left(\frac{r_{0}+1}{2}\right)^{q}\leq\frac{r^{q}_{0}+1}{2}.$
Applying the inequality for q=p1q=p-1 ($p\geq 2$) and multiplying both sides by $2^{p-1}$, gives us
$(r_{0}+1)^{p-1}\leq(r^{p-1}_{0}+1)2^{p-2}.$
Since $r^{p-1}_{0}+1=(r_{0}+1)(p-1)$, we get
$(r_{0}+1)^{p-1}\leq 2^{p-2}(r_{0}+1)(p-1).$
Dividing both sides by $r_{0}+1$, we obtain
$(r_{0}+1)^{p-2}\leq 2^{p-2}(p-1).$
Therefore, from (4.1), for $p\geq 2$, we have
$c_{1}(p)\geq\frac{1}{2^{p-2}}.$
By the Squeeze Theorem, it follows that $c_{1}(p)\to 0$ as $p\to\infty$. ∎
5. Proofs for the Lebesgue measure-Proofs of Theorem 3.3 and Corollary 3.6
Proof of Theorem 3.3.
Let $\Omega$ be a bounded convex domain of $\mathbb{R}^{n}$ ($n\geq 2$). Due to Theorem 3.1, we have that
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx=\int_{\Omega}C_{p}\left(\nabla u,u_{1}\nabla\left(\frac{u}{u_{1}}\right)\right)dx$
for all $u\in C_{0}^{\infty}(\Omega)$. Using Lemma 2.6, we have that for $p\geq 2$,
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq c_{1}(p)\int_{\Omega}\left|\nabla\left(\frac{u}{u_{1}}\right)\right|^{p}|u_{1}|^{p}dx.$ (5.1)
Let us define a function
$f:=\frac{u}{u_{1}}.$
The right hand side of (5.1) becomes
$c_{1}(p)\int_{\Omega}\left|\nabla f\right|^{p}|u_{1}|^{p}dx.$
Now, at this step, we would like to use Theorem 2.4. To do so, we need to show that $f$ is Lipschitz continuous. This follows from the regularity properties of $u_{1}>0$ and $u$ on bounded convex domains (see, Section 2). For convenience, let us denote $\omega=|u_{1}|^{p}$. Since Theorem 2.4 operates on log-concave measures, it remains to verify that $|u_{1}|^{p}$ is log-concave. By Theorem 2.1, the first eigenfunction satisfies that $\log u_{1}$ is concave. Observing that
$\log|u_{1}|^{p}=p\log u_{1},$
we conclude that $|u_{1}|^{p}$ is also log-concave. The last condition which needs to be satisfied is
$\int_{\Omega}|f|^{p-2}f\omega dx=0.$ (5.2)
However, this is not true in general. That is why we introduce a new function
$g(t):=\int_{\Omega}|f-t|^{p-2}(f-t)\omega dx.$
By the Dominated Convergence Theorem, we have the continuity of $g(t)$ with respect to $t$. Since $f$ is bounded, we have that
$m\leq f\leq M,$
where $m=\inf f$ and $M=\sup f$. Let us choose any $t<m$. Then,
$g(t)\geq(m-t)^{p-1}\int_{\Omega}\omega dx>0,$
which implies that $g$ attains a positive value. If $t>M$, then
$g(t)\leq-(t-M)^{p-1}\int_{\Omega}\omega dx<0,$
implying that $g$ also attains a negative value. Thus, since $g$ attains both negative and positive values and is continuous, by the Intermediate Value Theorem, there exists a constant $t_{0}$ such that $g(t_{0})=0$. This means
$\int_{\Omega}|f-t_{0}|^{p-2}(f-t_{0})\omega dx=0.$
Let us define
$\tilde{f}:=f-t_{0}.$
Since $\tilde{f}$ is Lipschitz continuous and it satisfies the condition (5.2), we can finally apply Theorem 2.4:
$c_{1}(p)\int_{\Omega}|\nabla\tilde{f}|^{p}|u_{1}|^{p}dx\geq c_{1}(p)C_{\Omega,p,\omega}\inf_{t\in\mathbb{R}}\int_{\Omega}|\tilde{f}-t|^{p}|u_{1}|^{p}dx.$
Making use of the identities, $\nabla\tilde{f}=\nabla f$ and $\tilde{f}-t:=f-c$, we get
$c_{1}(p)\int_{\Omega}|\nabla f|^{p}|u_{1}|^{p}dx\geq c_{1}(p)C_{\Omega,p,\omega}\inf_{c\in\mathbb{R}}\int_{\Omega}|f-c|^{p}|u_{1}|^{p}dx.$
Substituting back, we obtain
$c_{1}(p)\int_{\Omega}\left|\nabla\left(\frac{u}{u_{1}}\right)\right|^{p}|u_{1}|^{p}dx$ $\geq c_{1}(p)C_{\Omega,p,\omega}\inf_{c\in\mathbb{R}}\int_{\Omega}\left|\frac{u}{u_{1}}-c\right|^{p}|u_{1}|^{p}dx$
$=c_{1}(p)C_{\Omega,p,\omega}\inf_{c\in\mathbb{R}}\int_{\Omega}\left|u-cu_{1}\right|^{p}dx.$
In combination with (5.1), we get
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq c_{1}(p)C_{\Omega,p,\omega}\inf_{c\in\mathbb{R}}\int_{\Omega}\left|u-cu_{1}\right|^{p}dx.$
Finally, applying the bounds on the constants $c_{1}(p)$ and $C_{\Omega,p,\omega}$ from Proposition 2.10 and Theorem 2.4, respectively, we complete the proof. ∎
Proof of Corollary 3.6.
By Theorem 3.3, we have
$\int_{\Omega}|\nabla u|^{p}dx-\lambda_{1}(p,\Omega)\int_{\Omega}|u|^{p}dx\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}\inf_{c\in\mathbb{R}}\int_{\Omega}\left|u-cu_{1}\right|^{p}dx.$ (5.3)
We then extend the inequality (5.3) to $u\in W_{0}^{1,p}(\Omega)$ using the density argument. Now let $u=u_{2}$ be the second normalized eigenfunction of Dirichlet $p$-Laplacian, i.e. $\left\lVert u_{2}\right\rVert_{L^{p}(\Omega)}=1$ and
$\int_{\Omega}|\nabla u_{2}|^{p}dx=\lambda_{2}(p,\Omega)\int_{\Omega}|u_{2}|^{p}dx=\lambda_{2}.$
Then, in (5.3), we get
$\lambda_{2}(p,\Omega)-\lambda_{1}(p,\Omega)\geq\frac{1}{2^{p-2}}\left(\frac{\pi_{p}}{\text{diam}(\Omega)}\right)^{p}\inf_{c\in\mathbb{R}}\int_{\Omega}\left|u_{2}-cu_{1}\right|^{p}dx,$
which completes the proof. ∎
6. Proofs for the Gaussian probability measure–Proofs of Theorems 3.11 and 3.13
Proof of Theorem 3.11.
From Theorem 2.13, we have
$C_{p}(\xi,\eta)=R_{p}(\xi,\eta)=|\nabla u|^{p}-|\nabla\phi|^{p-2}\biggl\langle\nabla\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right),\nabla\phi\biggr\rangle.$
Multiplying both sides by $\left(2\pi\right)^{-\frac{n}{2}}e^{-\frac{|x|^{2}}{2}}$ and integrating over $\Omega$ give
$\int_{\Omega}C_{p}(\xi,\eta)d\gamma=\int_{\Omega}|\nabla u|^{p}d\gamma-\int_{\Omega}|\nabla\phi|^{p-2}\biggl\langle\nabla\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right),\nabla\phi\biggr\rangle d\gamma.$ (6.1)
To proceed next, we define the vector field
$X:=\frac{|u|^{p}}{|\phi|^{p-2}\phi}|\nabla\phi|^{p-2}\nabla\phi.$
Applying the product rule for the Gaussian divergence with $V=|\nabla\phi|^{p-2}\nabla\phi$ and $f=\frac{|u|^{p}}{|\phi|^{p-2}\phi}$, we get
$\text{div}_{\gamma}X=\biggl\langle\nabla\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right),\nabla\phi\biggr\rangle|\nabla\phi|^{p-2}+\frac{|u|^{p}}{|\phi|^{p-2}\phi}\operatorname{div}_{\gamma}\left(|\nabla\phi|^{p-2}\nabla\phi\right).$ (6.2)
Note that by definition $\text{div}_{\gamma}\left(|\nabla\phi|^{p-2}\nabla\phi\right)=\Delta_{p,\gamma}\phi$. Rewriting (6.2):
$|\nabla\phi|^{p-2}\biggl\langle\nabla\left(\frac{|u|^{p}}{|\phi|^{p-2}\phi}\right),\nabla\phi\biggr\rangle=\text{div}_{\gamma}X-\frac{|u|^{p}}{|\phi|^{p-2}\phi}\Delta_{p,\gamma}\phi.$ (6.3)
Substituting (6.3) into (6.1):
$\int_{\Omega}C_{p}\left(\xi,\eta\right)d\gamma=\int_{\Omega}|\nabla u|^{p}d\gamma-\int_{\Omega}\text{div}_{\gamma}Xd\gamma+\int_{\Omega}\frac{|u|^{p}}{|\phi|^{p-2}\phi}\Delta_{p,\gamma}\phi d\gamma.$
After integrating by parts and substituting back $\xi$ and $\eta$, we get
$\int_{\Omega}|\nabla u|^{p}d\gamma+\int_{\Omega}\frac{|u|^{p}}{|\phi|^{p-2}\phi}\Delta_{p,\gamma}\phi d\gamma=\int_{\Omega}C_{p}\left(\nabla u,\phi\nabla\left(\frac{u}{\phi}\right)\right)d\gamma.$
Finally, setting $\phi$ to be the first eigenfunction $u^{\gamma}_{1}$ of the Gaussian $p$-Laplacian, we obtain
$\int_{\Omega}|\nabla u|^{p}d\gamma-\lambda_{1}(p,\Omega,\gamma)\int_{\Omega}|u|^{p}d\gamma=\int_{\Omega}C_{p}\left(\nabla u,u_{1}^{\gamma}\nabla\left(\frac{u}{u_{1}^{\gamma}}\right)\right)d\gamma.$
The proof is complete. ∎
Proof of Theorem 3.13.
Set $\Omega$ to be a bounded convex domain of $\mathbb{R}^{n}$ ($n\geq 2$). By Theorem 3.6, we have
$\int_{\Omega}|\nabla u|^{p}d\gamma-\lambda_{1}(p,\Omega,\gamma)\int_{\Omega}|u|^{p}d\gamma=\int_{\Omega}C_{p}\left(\nabla u,u^{\gamma}_{1}\nabla\left(\frac{u}{u_{1}}\right)\right)d\gamma$
for all real-valued $u\in C_{0}^{\infty}(\Omega)$. Lemma 2.6 implies that, for $p\geq 2$, we have
$\int_{\Omega}|\nabla u|^{p}d\gamma-\lambda_{1}(p,\Omega,\gamma)\int_{\Omega}|u|^{p}d\gamma\geq c_{1}(p)\int_{\Omega}\left|\nabla\left(\frac{u}{u_{1}^{\gamma}}\right)\right|^{p}|u^{\gamma}_{1}|^{p}d\gamma.$ (6.4)
Now we define
$f:=\frac{u}{u^{\gamma}_{1}}.$
The regularity properties of $u$ and $u^{\gamma}_{1}$ imply that $f$ is Lipschitz continuous (see, Proposition 2.3). Since $|u^{\gamma}_{1}|^{p}$ is log-concave by Theorem 2.2, we have that
$|u^{\gamma}_{1}|^{p}\left(2\pi\right)^{-\frac{n}{2}}e^{-\frac{|x|^{2}}{2}}$
is also log-concave. Let us denote $\omega=|u^{\gamma}_{1}|^{p}\left(2\pi\right)^{-\frac{n}{2}}e^{-\frac{|x|^{2}}{2}}$. Then, by an analogous argument from the proof of Theorem 3.3, we then eventually get from Theorem 2.4:
$c_{1}(p)\int_{\Omega}|\nabla f|^{p}|u^{\gamma}_{1}|^{p}d\gamma\geq C_{\Omega,p,\omega}\inf_{c\in\mathbb{R}}\int_{\Omega}|f-c|^{p}|u^{\gamma}_{1}|^{p}d\gamma$
for $C_{\Omega,p,\omega}>0$. After the change of variables and applying it in combination with the estimate (6.4), we get
$\int_{\Omega}|\nabla u|^{p}d\gamma-\lambda_{1}(p,\Omega,\gamma)\int_{\Omega}|u|^{p}d\gamma\geq c_{1}(p)C_{\Omega,p,\omega}\inf_{c\in\mathbb{R}}\int_{\Omega}|u-cu^{\gamma}_{1}|^{p}d\gamma.$
Applying the lower bounds on the constants $c_{1}(p)$ and $C_{\Omega,p,\omega}$ from Proposition 2.10 and Theorem 2.4, respectively, we complete the proof. ∎
Acknowledgments. We would like to thank Vladimir Bobkov for valuable discussions and comments as well as for pointing out an important reference [FT02] in the field.